OpenStudy (anonymous):

Solve the following system of equations using elimination: 3x - 4y = 8 10x + 5y = 45 A. (-4, -1) B. (3, 3) C. (4, 1) D. (8, 4)

OpenStudy (anonymous):

3x - 4y = 8 5(3x - 4y = 8) 10x + 5y = 45 4(10x + 5y = 45)

OpenStudy (anonymous):

15x - 20y=40 40x+20y=180 55x = 220 x=4 3(4)-4y=8 12-4y=8 -4y=-4 y=1 solution is (4,1) check 2nd equation: 10x + 5y = 45 10(4)+5(1)=45 40+5 = 45 45=45 solution works